# ECG-Chat: A Large ECG-Language Model for Cardiac Disease Diagnosis

### v1

@@ Introduction
ECG-Chat is trained in two stages. The training data for both are obtained from MIMIC-IV-ECG (Gow et al. 2023). The pretraining dataset includes a 619K records, each record has a random GPT-4 generated user instruction, and the reports in the dataset constitute the LLM’s answer. The fine-tuning dataset includes 19K diagnosis data and 25K conversation data. The diagnosis data is similar to the pretraining data, but the answer template is more in line with user interaction. The conversation data set contains 4-15 rounds of dialogue about ECG details. The context and training data when constructing the fine-tuning dataset are shown in Figure 4. The specific construction process is shown in Appendix E.
Datasets for Evaluations
@@
For ECG encoder, we evaluate on ECG-Report retrieval and zero-shot, linear probe classification. For ECG-Chat, we evaluate on ECG report generation task. The datasets used for evaluation are PTB-XL (Wagner et al. 2020) and CPSC2018 (Liu et al. 2018) datasets. For the ablation study of GraphRAG and DSPy, we use the ECG-ExpertQA dataset, which is constructed by chatGPT-4o. More information can be found in Appendix D.
Experiment
Implementation Details
@@
In the ECG encoder training phase, we used a 12-layer 1d-ViT (Dosovitskiy et al. 2020) as the backbone. ECG-ViT has a patch size of 50, a hidden size of 768, 12 heads, and an MLP size of 3072. For the text encoder, we used the model Med-CPT (Jin et al. 2023) obtained on the text contrastive learning task and only fine-tuned its last two layers. The embedding dimension of the dual encoder is 512. The text decoder is also a 6-layer transformer decoder (Vaswani et al. 2017). During training, the batch size per GPU is 128. The AdamW optimizer with a learning rate of 1e-4 and a weight decay of 0.1 is used for 20 epochs. The weights of contrastive loss and captioning loss are 1.0 and 2.0 respectively. In addition, we used three ECG data augmentation strategies: baseline wander, cut mix and random masking (Wen and Kang 2022).
@@
In ECG-Chat, the projection layer is a two-layer MLP, and the LLM base is Vicuna-13B (Zheng et al. 2023). AdamW optimizer with a cosine learning rate scheduler is used to train our model. The pretraining stage lasts for 1 epoch, where only the projection layer is trained. The fine-tuning stage lasts for 3 epochs, where the projection layer is trained and Vicuna-13B is fine-tuned with LoRA (Hu et al. 2021).In order to save memory, ZeRO (Rajbhandari et al. 2020) was used in the construction of ECG-Chat.
All experiments were run on 8×V100 32GB GPUs using PyTorch.
Zero-shot ECG-Report Retrieval
@@
Table 1: Zero-shot ECG-report retrieval results on PTB-XL (2K test set)
Model PTB-XL Super PTB-XL Sub PTB-XL Form PTB-XL Rhythm CPSC 2018
MoCo-v3 64.87/84.72 32.15/83.21 27.83/71.29 50.65/89.13 70.45/90.28
Simsiam 66.35/85.94 34.89/84.67 30.64/73.58 52.48/90.52 71.93/91.47
CRT 68.02/86.22 40.22/87.45 33.21/75.11 53.79/91.02 73.26/92.12
MERL 67.13/86.50 36.56/85.05 29.47/74.49 48.12/91.99 72.80/92.62
CoCa 73.91/91.42 50.87/89.52 39.61/85.19 58.10/93.80 80.28/95.62
CoCa+WDE 72.20/90.59 45.10/89.34 38.10/82.05 55.92/94.04 80.10/95.72
Table 2: Performance comparison of different models across various datasets, with combined F1-Score and AUC values.
@@
Table 3: Ablation study of DDP in clinical efficacy (CE) and natural language generation (NLG) metrics evaluation
Table 1 shows the retrieval results of ECG-Chat’s ECG encoder on the PTB-XL test set of size 2K, and compares it with models that do not use Waveform Data Enhancement (WDE), including All-Grid (Qiu et al. 2023a) which using image encoding, ALBEF (Li et al. 2021), MERL (Liu et al. 2024b), and CoCa (Yu et al. 2022) models. ALBEF, MERL, and CoCa use the same ECG ViT encoder, pretrained text encoder and training datasets as our model. As can be seen from Table 1, the CoCa model achieves the best results in retrieval. Also, without WDE, the recall of the retrieval is very low. WDE increases the differences between ECG reports and provides more information. The CoCa model with WDE achieves the best results.
ECG Classification
ECG classification is crucial for intelligent heart disease diagnosis, aiming to develop and evaluate a model that accurately identifies arrhythmias and other cardiac conditions. This experiment tests the ECG ViT encoder’s feature extraction capabilities and its generalization and adaptability in transfer scenarios. We used five datasets, including four PTB-XL subsets (Super, Sub, Form, Rhythm) and the CPSC 2018 dataset, covering a broad range of cardiac conditions to challenge the model’s generalization. Performance was evaluated using F1-Score and AUC.
Table 2 shows the effect of our ECG encoder on the linear probing classification task, and compares it with the current state-of-the-art (SOTA) model MERL (Liu et al. 2024b) and three ECG self-supervised learning (eSSL) methods, including MoCo-v3 (Chen, Xie, and He 2021), Simsiam (Chen and He 2021), and CRT (Zhang et al. 2023b). The CoCa model achieved the best results in F1 and AUC on the five datasets. After adding WaveForm Data, although the scores decreased slightly in some datasets, the results still performed very well. However, we also observed that all models performed poorly on the Form dataset, even with WDE. This is also an area that needs improvement.
ECG Reports Generation
@@
Figure 5: ECG-Chat F1 score statistics for different categories. The left side is the Disease subset, and the right side is the Form and Rhythm subsets. The labels are abbreviated as SCP Codes(Rubel et al. 2016).
GraphRAG DSPy F AR CR CP CU CER SS
39.87 32.56 9.03 29.82 29.13 24.67 18.94
76.60 68.29 32.67 67.53 64.39 57.00 54.40
71.57 63.93 27.57 62.81 59.54 53.35 47.72
82.12 74.29 39.44 73.18 88.92 73.10 81.83
Table 4: Ablation Study of GraphRAG and DSPy in Model Performance Metrics
For evaluation, both clinical efficacy (CE) and natural language generation (NLG) metrics are used. CE includes Precision, Recall and F1 score, where the labels of text reports are annotated by GPT-4. NLG includes BLEU (Papineni et al. 2002), METEOR (Denkowski and Lavie 2011) and ROUGE-L (Lin 2004). NLG meteics uses GPT-4o’s answers based on PTB-XL translated reports as reference.
Table 3 shows the NLG evaluation results of ECG-Chat. We compare it with the translated original report in PTB-XL. ECG-Chat with DDP achieves the best results in all indicators. The original report in PTB-XL is shorter, so it is not as good as ECG-Chat in BLUE and METEOR metrics. At the same time, ECG-Chat is also better than the original report in ROUGE-1 and ROUGE-L, which may be because ECG-Chat has the same template as the reference GPT-4o response, and DDP also greatly improves the accuracy of the answer.
Table 3 also shows the results of DDP on the CE metrics. Due to the small number of training data sets, the model without DDP showed serious hallucinations. Therefore, at the current stage, this prompt is indispensable. However, on the three data sets, the recall is relatively low, indicating that the model has some problems in judging negative samples.
@@
Effectiveness of GraphRAG and DSPy
Table 4 shows the combination of GraphRAG and DSPy significantly outperforms other combinations across all metrics. Without GraphRAG and DSPy, the model’s performance is relatively poor, with a Faithfulness score of 39.87 and Answer Relevancy score of 32.56, among others. Introducing the GraphRAG module results in a significant improvement in Faithfulness (76.60) and Answer Relevancy (68.29). Meanwhile, incorporating the DSPy module leads to improvements in Context Recall (27.57) and Summarization Score (47.72).
Ablation Study
In this section, we study the performance of different modules from three aspects: dataset, text encoder, scalability, and evaluate the impact on the ECG encoder on retrieval and classification tasks. The specific experimental settings and results are shown in Appendix F.
Conclusion and Limitations
@@
Although our model has achieved excellent results on multiple tasks, there are still some limitations. First, due to the lack of diversity in the training dataset, the ECG features cannot be aligned with the LLM model as expected. At the same time, the dataset for instruction tuning is small and does not come from real world, which leads to bias and hallucinations in LLMs. Second, the model can still be improved in diagnosis, especially in waveforms and some rare symptoms. Finally, our model only focuses on ECG, and we also hope to combine it with the most advanced medical large language model (Chen et al. 2024b; Tu et al. 2024), integrating multiple modalities such as X-Ray, EHR, etc., to provide a more comprehensive interpretation of medical records.
Appendix A Appendix A: RAG of Cardiology Knowledge
@@
ECG-Chat uses Microsoft’s GraphRAG (Edge et al. 2024) component to convert the content from professional books such as ”ECG Workout - Exercises in Arrhythmia Interpretation (Huff 2022)”, ”Manual of Cardiovascular Medicine (Griffin 2018)”, ”Medical Student Survival Skills ECG (Jevon and Gupta 2019)”, ”Cardiology Subspecialty Consult (Crawford and Lin 2004)”, ”The ECG Made Easy (Hampton and Hampton 2019)”, ”The ECG In Practice (Hampton 2017)”, and ”Arrhythmia Recognition: The Art of Interpretation (Miller and Garcia 2003)” into a graph index. This index, built through a knowledge graph of nodes and edges, comprehensively covers the knowledge of ECG interpretation and cardiovascular diagnosis. GraphRAG first parses the content of these books into a graph structure and uses community detection algorithms to group related medical topics. When a user submits a query, the system retrieves and summarizes relevant elements from the graph index, generating a comprehensive ”global answer,” thereby helping to mitigate hallucinations in LLMs during the generation of ECG diagnostic reports. Ablation experiments have validated that this module is effective.
Appendix B Appendix B: Automated Prompt Tuning
@@
We employ a DSPy component (Soylu, Potts, and Khattab 2024), trained on the ECG-Instruct dataset, to perform automatic prompt tuning. DSPy leverages GraphRAG to retrieve relevant knowledge such as symptom causes, medication information, and clinical guidelines, and combines this knowledge with the patient’s vital signs and cardiology diagnostic data to form a comprehensive medical context. Based on this, the DSPy module automatically generates accurate and detailed medical diagnostic texts through a series of declarative programming steps. The automation and modular design of this workflow enable DSPy to optimize the prompts and weights of the language model, improving the accuracy and efficiency of the output while reducing the need for manual tuning, thereby enhancing the system’s scalability and maintainability.
Appendix C Appendix C: LaTeX-based Automated Medical Report Generation Pipeline
@@
Clinical diagnosis aims for structured output to enhance the evaluation and readability of reports. ECG-Chat employs plasTeX (Smith 2023) to generate clinical diagnostic reports in a highly structured and automated manner. Initially, patient personal information, including basic details and key medical background, is collected and integrated from outpatient systems or patient databases, forming the ”Patient Information” and ”Medical History” sections. Subsequently, ECG-Chat analyzes ECG data to extract key metrics and patterns, which are detailed in the ”ECG Data Analysis” section. The report then incorporates pathological analysis results, providing thorough descriptions and interpretations in the ”Pathological Analysis” section. Based on the consolidated data and analyses, ECG-Chat generates the ”Diagnosis” section, which explicitly states the clinical diagnosis. Finally, the system proposes specific treatment recommendations and preventive measures, which are outlined in the ”Recommendations” section. A sample of this template is illustrated in Figure 7 and Figure 7.
Appendix D Appendix D: Description of Datasets
@@
  • MIMIC-IV-ECG: This dataset contains 800,035 ECG records across nearly 160,000 unique patients (Gow et al. 2023). Every ECG recording sampled at 500Hz for a duration of 10 seconds and Each ECG corresponds to several machine-generated text reports. We process this dataset using the following strategy: (1) Replace ’NAN’ and ’inf’ with 0 in ECG data. (2) Text in the reports that are not relevant to the diagnosis are excluded, and samples whose final report is empty are removed. Finally, we use 788,822 samples as training data.
  • Champan-Shaoxing-Ningbo (CSN): This dataset contains 45,152 12-lead ECG records (Zheng et al. 2020; Zheng, Guo, and Chu 2022). Every ECG recording is also sampled at 500Hz for a duration of 10 seconds. Each record is annotated by several SNOMED CT codes111 https://www.snomed.org/. We converted each SNOMED CT code into a corresponding textual description, merged as a report of the ECG record, and used it for data training. The number of training samples is 40,637.
  • Shandong Provincial Hospital (SPH) database: The dataset contains 25,770 ECG records from 24,666 patients (Liu et al. 2022). The length of each recording is between 10 and 60 seconds, and the sampling rate is 500Hz. We intercept the first 10 seconds in the record as training data. Each recorded diagnosis in the dataset corresponds to standardized diagnostic statements conforming to the AHA/ACC/HRS recommendations (Kligfield et al. 2007). Similarly, we also convert these statements into reports of corresponding records. The number of training samples is 20,616.
  • PTB-XL: This dataset contains 21,837 10s ECG signals that come from 18,885 patients. Each ECG record has a corresponding free text report and SCP Codes label (Rubel et al. 2016) extracted from the report.
@@
      Conversations: This dataset was designed by GPT-4o for multi-turn ECG conversations between patients and physicians, including heart rate, waveform, rhythm, cardiac axis, and diagnostic results. Therefore, in addition to the text reports, the calculated waveform data by NeuroKit2 (Makowski et al. 2021) on lead II were included in the prompt we provided to GPT-4o. Additionally, in order to enhance the robustness of the model, diagnostic or waveform features that did not exist in the original reports were randomly added to the prompt words, leading to negative answers in the data, as context type 3 shown in Fig. 4. This dataset contains 25K samples, each sample has 4-15 rounds dialogue. Of the 25K samples, 19K are from the diagnosis dataset.
@@
Appendix E Appendix E: Dataset Statistics and Analysis on ECG-Instruct
@@
Table 5: Statistical indicators of the ECG-Instruct dataset.
@@
Table 5 and Figure 8 present the statistical indicators of the ECG-Instruct dataset. The dataset contains 45,044 ECG samples with a total of 325,730 sentences, averaging 7.23 sentences per description. Each description has an average length of 73.15 vocabularies, amounting to a total of 3,294,880 vocabularies, including 6,319 distinct vocabularies. These statistics indicate that the ECG-Instruct dataset exhibits a high level of granularity in its textual descriptions, offering detailed explanations of ECG signals and their clinical implications. This rich textual information provides a solid foundation for understanding and automatically generating clinical diagnostic reports.
@@
Appendix F Appendix F: Ablation Study
Dataets. During the ECG encoder training process, we merged three training datasets. Table 6 compares the effects of a single MIMIC-IV-ECG dataset and a mixed dataset. It can be found that the mixed dataset can effectively enhance the generalization ability of the model.

### v2

@@ IV-B Datasets for Instrution Tuning
ECG-Chat is trained in two stages. The training data for both are obtained from MIMIC-IV-ECG [27]. The pretraining dataset includes a 619K records, each record has a random GPT-4 generated user instruction, and the reports in the dataset constitute the LLM’s answer. The fine-tuning dataset includes 19K diagnosis data and 25K conversation data. The diagnosis data is similar to the pretraining data, but the answer template is more in line with user interaction. The conversation data set contains 4-15 rounds of dialogue about ECG details. The context and training data when constructing the fine-tuning dataset are shown in Figure 3. The specific construction process is shown in Appendix E.
IV-C Datasets for Evaluations
@@
For ECG encoder, we evaluate on ECG-Report retrieval and zero-shot, linear probe classification. For ECG-Chat, we evaluate on ECG report generation task. The datasets used for evaluation are PTB-XL [32] and CPSC2018 [33] datasets. For the ablation study of GraphRAG and DSPy, we use the ECG-ExpertQA dataset, which is constructed by chatGPT-4o. More information can be found in Appendix D.
V Experiment
V-A Implementation Details
@@
In the ECG encoder training phase, we used a 12-layer 1d-ViT as the backbone. ECG-ViT has a patch size of 50, a hidden size of 768, 12 heads, and an MLP size of 3072. For the text encoder, we used the model Med-CPT [34] obtained on the text contrastive learning task and only fine-tuned its last two layers. The embedding dimension of the dual encoder is 512. The text decoder is also a 6-layer transformer decoder. During training, the batch size per GPU is 128. The AdamW optimizer with a learning rate of 1e-4 and a weight decay of 0.1 is used for 20 epochs. The weights of contrastive loss and captioning loss are 1.0 and 2.0 respectively. In addition, we used three ECG data augmentation strategies: baseline wander, cut mix and random masking [35].
@@
In ECG-Chat, the projection layer is a two-layer MLP, and the LLM base is Vicuna-13B [7]. AdamW optimizer with a cosine learning rate scheduler is used to train our model. The pretraining stage lasts for 1 epoch, where only the projection layer is trained. The fine-tuning stage lasts for 3 epochs, where the projection layer is trained and Vicuna-13B is fine-tuned with LoRA [24].In order to save memory, ZeRO [36] was used in the construction of ECG-Chat.
All experiments were run on 8×V100 32GB GPUs.
V-B Zero-shot ECG-Report Retrieval
@@
TABLE I: Zero-shot ECG-report retrieval results on PTB-XL (2K test set)
@@
TABLE II: Ablation study of DDP in clinical efficacy (CE) and natural language generation (NLG) metrics evaluation
@@
Table I shows the retrieval results of ECG-Chat’s ECG encoder on the PTB-XL test set of size 2K, and compares it with models that do not use Waveform Data Enhancement (WDE), including All-Grid [37] which using image encoding, ALBEF [38], MERL [5], and CoCa [19] models. ALBEF, MERL, and CoCa use the same ECG ViT encoder, pretrained text encoder and training datasets as our model. As can be seen from Table I, the CoCa model achieves the best results in retrieval. Also, without WDE, the recall of the retrieval is very low. WDE increases the differences between ECG reports and provides more information. The CoCa model with WDE achieves the best results.
V-C ECG Reports Generation
@@
Figure 4: ECG-Chat F1 score statistics for different categories. The left side is the Disease subset, and the right side is the Form and Rhythm subsets. The labels are abbreviated as SCP Codes[39].
For evaluation, both clinical efficacy (CE) and natural language generation (NLG) metrics are used. CE includes Precision, Recall and F1 score, where the labels of text reports are annotated by GPT-4. NLG includes BLEU, METEOR and ROUGE-L. NLG meteics uses GPT-4o’s answers based on PTB-XL translated reports as reference.
Table II shows the NLG evaluation results of ECG-Chat. We compare it with the translated original report in PTB-XL. ECG-Chat with DDP achieves the best results in all indicators. The original report in PTB-XL is shorter, so it is not as good as ECG-Chat in BLUE and METEOR metrics. At the same time, ECG-Chat is also better than the original report in ROUGE-1 and ROUGE-L, which may be because ECG-Chat has the same template as the reference GPT-4o response, and DDP also greatly improves the accuracy of the answer.
Table II also shows the results of DDP on the CE metrics. Due to the small number of training data sets, the model without DDP showed serious hallucinations. Therefore, at the current stage, this prompt is indispensable. However, on the three data sets, the recall is relatively low, indicating that the model has some problems in judging negative samples.
@@
VI Conclusion and Limitations
@@
Although our model has achieved excellent results on multiple tasks, there are still some limitations. First, due to the lack of diversity in the training dataset, the ECG features cannot be aligned with the LLM model as expected. At the same time, the dataset for instruction tuning is small and does not come from real world, which leads to bias and hallucinations in LLMs. Second, the model can still be improved in diagnosis, especially in waveforms and some rare symptoms. Finally, our model only focuses on ECG, and we also hope to combine it with the most advanced medical large language model [40], integrating multiple modalities such as X-Ray, EHR, etc., to provide a more comprehensive interpretation of medical records.
Appendix A RAG of Cardiology Knowledge
@@
ECG-Chat uses Microsoft’s GraphRAG [41] component to convert the content from professional books such as ”ECG Workout - Exercises in Arrhythmia Interpretation [42]”, ”Manual of Cardiovascular Medicine [43]”, ”Medical Student Survival Skills ECG [44]”, ”Cardiology Subspecialty Consult [45]”, ”The ECG Made Easy [46]”, ”The ECG In Practice [47]”, and ”Arrhythmia Recognition: The Art of Interpretation [48]” into a graph index. This index, built through a knowledge graph of nodes and edges, comprehensively covers the knowledge of ECG interpretation and cardiovascular diagnosis. GraphRAG first parses the content of these books into a graph structure and uses community detection algorithms to group related medical topics. When a user submits a query, the system retrieves and summarizes relevant elements from the graph index, generating a comprehensive ”global answer,” thereby helping to mitigate hallucinations in LLMs during the generation of ECG diagnostic reports. Ablation experiments have validated that this module is effective.
Appendix B Automated Prompt Tuning
@@
We employ a DSPy component [49], trained on the ECG-Instruct dataset, to perform automatic prompt tuning. DSPy leverages GraphRAG to retrieve relevant knowledge such as symptom causes, medication information, and clinical guidelines, and combines this knowledge with the patient’s vital signs and cardiology diagnostic data to form a comprehensive medical context. Based on this, the DSPy module automatically generates accurate and detailed medical diagnostic texts through a series of declarative programming steps. The automation and modular design of this workflow enable DSPy to optimize the prompts and weights of the language model, improving the accuracy and efficiency of the output while reducing the need for manual tuning, thereby enhancing the system’s scalability and maintainability.
Appendix C LaTeX-based Automated Medical Report Generation Pipeline
@@
Clinical diagnosis aims for structured output to enhance the evaluation and readability of reports. ECG-Chat employs plasTeX [50] to generate clinical diagnostic reports in a highly structured and automated manner. Initially, patient personal information, including basic details and key medical background, is collected and integrated from outpatient systems or patient databases, forming the ”Patient Information” and ”Medical History” sections. Subsequently, ECG-Chat analyzes ECG data to extract key metrics and patterns, which are detailed in the ”ECG Data Analysis” section. The report then incorporates pathological analysis results, providing thorough descriptions and interpretations in the ”Pathological Analysis” section. Based on the consolidated data and analyses, ECG-Chat generates the ”Diagnosis” section, which explicitly states the clinical diagnosis. Finally, the system proposes specific treatment recommendations and preventive measures, which are outlined in the ”Recommendations” section. A sample of this template is illustrated in Figure 6 and Figure 6.
Appendix D Description of Datasets
@@
  • MIMIC-IV-ECG: This dataset contains 800,035 ECG records across nearly 160,000 unique patients [27]. Every ECG recording sampled at 500Hz for a duration of 10 seconds and Each ECG corresponds to several machine-generated text reports. We process this dataset using the following strategy: (1) Replace ’NAN’ and ’inf’ with 0 in ECG data. (2) Text in the reports that are not relevant to the diagnosis are excluded, and samples whose final report is empty are removed. Finally, we use 788,822 samples as training data.
  • Champan-Shaoxing-Ningbo (CSN): This dataset contains 45,152 12-lead ECG records [28, 29]. Every ECG recording is also sampled at 500Hz for a duration of 10 seconds. Each record is annotated by several SNOMED CT codes111 https://www.snomed.org/. We converted each SNOMED CT code into a corresponding textual description, merged as a report of the ECG record, and used it for data training. The number of training samples is 40,637.
  • Shandong Provincial Hospital (SPH) database: The dataset contains 25,770 ECG records from 24,666 patients [30]. The length of each recording is between 10 and 60 seconds, and the sampling rate is 500Hz. We intercept the first 10 seconds in the record as training data. Each recorded diagnosis in the dataset corresponds to standardized diagnostic statements conforming to the AHA/ACC/HRS recommendations [51]. Similarly, we also convert these statements into reports of corresponding records. The number of training samples is 20,616.
  • PTB-XL: This dataset contains 21,837 10s ECG signals that come from 18,885 patients. Each ECG record has a corresponding free text report and SCP Codes label [39] extracted from the report.
@@
      Conversations: This dataset was designed by GPT-4o for multi-turn ECG conversations between patients and physicians, including heart rate, waveform, rhythm, cardiac axis, and diagnostic results. Therefore, in addition to the text reports, the calculated waveform data by NeuroKit2 [31] on lead II were included in the prompt we provided to GPT-4o. Additionally, in order to enhance the robustness of the model, diagnostic or waveform features that did not exist in the original reports were randomly added to the prompt words, leading to negative answers in the data, as context type 3 shown in Fig. 3. This dataset contains 25K samples, each sample has 4-15 rounds dialogue. Of the 25K samples, 19K are from the diagnosis dataset.
@@
Appendix E Dataset Statistics and Analysis on ECG-Instruct
@@
TABLE III: Statistical indicators of the ECG-Instruct dataset.
@@
Table III and Figure 7 present the statistical indicators of the ECG-Instruct dataset. The dataset contains 45,044 ECG samples with a total of 325,730 sentences, averaging 7.23 sentences per description. Each description has an average length of 73.15 vocabularies, amounting to a total of 3,294,880 vocabularies, including 6,319 distinct vocabularies. These statistics indicate that the ECG-Instruct dataset exhibits a high level of granularity in its textual descriptions, offering detailed explanations of ECG signals and their clinical implications. This rich textual information provides a solid foundation for understanding and automatically generating clinical diagnostic reports.
@@
Appendix F Ablation Study
Dataets
During the ECG encoder training process, we merged three training datasets. Table IV compares the effects of a single MIMIC-IV-ECG dataset and a mixed dataset. It can be found that the mixed dataset can effectively enhance the generalization ability of the model.